OpenStudy (anonymous):

A meter stick flies by an apparent length of 60 cm. What is its velocity (in m/s)?

OpenStudy (amistre64):

this sounds like a relativity question dealing with contraction of space

OpenStudy (anonymous):

\[l=l _{0}\sqrt{1-\frac{ v ^{2} }{ c ^{2} }}\]\[\frac{ 60 }{ 100}=0.6=\sqrt{1-\frac{ v ^{2} }{ c ^{2} }}\] as, \[\frac{ l }{ l _{0} }=\sqrt{1-\frac{ v ^{2} }{ c^{2} }}\] therefore, squaring both the side, \[0.36=1-\frac{ v ^{2} }{ c^{2} }\] \[\frac{ v ^{2} }{ c ^{2} }=0.64\] taking square root,\[v=0.8c\]